mrs. williams mixed nuts that cost $3.90 per pound with nuts that cost $4.30 per pound to obtain a mixture of 50 pounds of nuts worth $4.20 per pound. how many pounds of each type of nuts did she use?
Answer by CharlesG2(834)

mrs. williams mixed nuts that cost $3.90 per pound with nuts that cost $4.30 per pound to obtain a mixture of 50 pounds of nuts worth $4.20 per pound. how many pounds of each type of nuts did she use?

price/lb A * amount A + price/lb B * amount B = price/lb C * amount C
$3.90/lb * A + $4.30/lb * B = $4.20/lb * C
let amount C = amount A + amount B = 50 lbs, let amount B = 50 - A lbs
3.90A + 4.30(50 - A) = 4.20(50)
3.90A + 215 - 4.30A = 210
-0.40A = -5
0.40A = 5 (divided the -1 out)
A = 12.5 lbs of one type of nuts
B = 50 - A = 50 - 12.5 = 37.5 lbs of the other type of nut
we could tell before solving anything that there was going to be more of the $4.30 nuts than the $3.90 nuts since $4.30 is closer to $4.20 than $3.90 is